### v1

# Contrastive Chain-of-Thought Prompting

Abstract
Rapidly increasing model scales coupled with steering methods such as chain-of-thought prompting (Wei et al., 2022b) have led to drastic improvements in language model reasoning. At the same time, models struggle with compositional generalization and are far from human performance on many reasoning-based benchmarks. Leveraging the success of chain-of-thought prompting, and also taking inspiration from context-aware decoding (CAD) (Shi et al., 2023), we explore input-based contrasting methods to further encourage the type of reasoning induced by chain-of-thought prompting. While work remains to stabilize these results across datasets and models, the improvements we find warrant further investigation into input-based steering methods for context-aware reasoning.
Contrastive Chain-of-Thought Prompting
Grant Kruttschnitt           Jay Shim           Alyssa Ma           Daniel Kim Benjamin Chek           Athul Anand                     Kevin Zhu           Sean O’Brien Algoverse AI Research sean@algoverse.us, kevin@algoverseacademy.com
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing by leveraging vast amounts of data to generate human-like text. These models have gained attention for their ability to perform a wide range of language tasks, from text generation and translation to question answering and summarization. One such technique that has propelled the performance of LLMs is Chain-of-Thought (CoT) prompting (Wei et al., 2022b) . By applying CoT prompting, researchers have found substantial improvement in various tasks such as arithmetic and common sense reasoning, surpassing the efficacy of traditional methods focused on scaling model size (Wang et al., 2023; Wei et al., 2023). In a similar vein, Context-Aware Decoding (CAD) emerges as another approach to augment LLM performance (Liu et al., 2021; Li et al., 2023a; O’Brien and Lewis, 2023), priotizing relevant contextual information over pre-existing knowledge (Shi et al., 2023). While these advancements underscore significant progress, an ongoing question persists regarding the efficacy of combining CAD with CoT or with amateur prompts, which typically contain less context.
To address this question, the study aims to investigate how CoT prompts’ ability to capture sequential dependencies can be complemented by the discriminative ability of CAD, particularly when the context contradicts prior knowledge (Longpre et al., 2021; Zhou et al., 2023). In such cases, we aim to ensure that incorporating CoT prompting does not detract from the outputs derived from the provided context. Ultimately, the goal is to create faithful models capable of accurately executing mathematical and reasoning tasks while avoiding the misinterpretation of crucial contextual cues (Maynez et al., 2020; Pagnoni et al., 2021).
2 Method
2.1 Background
We take a query x which we feed to our language model $\theta$, to generate a response y using greedy decoding. The response is sampled autoregressively from two probability distributions given by query $x$, CoT prompt c, and an amateur model (3.2.1) with less context amt (Shi et al., 2023). Using a contrastive strength of $\alpha$, we create a new distribution:
$y_{t}\propto\frac{p_{\theta}(y_{t}|\mathbf{c,x,y_{<t}})^{\alpha+1}}{p_{\theta}(y_{t}|\mathbf{amt,x,y_{<t}})^{\alpha}}$
We hypothesize that the model can effectively rely on the prompt c when reasoning with the input while simultaneously penalizing incorrect reasoning behaviors and conciseness (amt) to provide an accurate output.
2.2 Context-Aware Decoding
For each method we utilize two different prompts: an 8-shot CoT prompt, and an amateur prompt. To identify the next logit and thus the next token to add autoregressively, we use the contrastive function below proposed by Shi et al. (2023):
$y_{t}\sim\textrm{softmax}\left[(1+\alpha)\exp(\text{logit}_{\theta}(y_{t}|\mathbf{c,x,y_{<t}}))\right.$
$\left.-\alpha\exp(\text{logit}_{\theta}(y_{t}|\mathbf{amt,x,y_{<t}}))\right]$
The values of $\alpha$ are as follows, $0\leq\alpha\leq 1$, with higher alpha values indicating stronger reliance on prompt c, the CoT prompt. From the adjusted logits, we apply greedy decoding to generate the subsequent character. In our analysis, a single model was used to generate probabilities for both CoT and amateur prompts, but different models may also be used.
3 Experimental Setup
3.1 Datasets
We evaluated our method on three question-answering task datasets that measure performance based on reasoning abilities: GSM8K (Cobbe et al., 2021), AQuA (Ling et al., 2017), and CommonSenseQA (Talmor et al., 2019). We provide multiple-choice options in the prompts of CommonSenseQA and AQuA, but do not for GSM8K. GSM8K and AQuA are math datasets, while CommonSenseQA contains common-sense reasoning problems.
3.2 Models and Prompting
We performed our experiments using Phi-1.5, Mistral 7B, and GPT-3.5 with version 3.5-turbo-0125. In each of our experiments, we contrasted the logits obtained from a 8-shot CoT prompt with the amateur. The exemplars in the few-shot prompts are taken from the corresponding dataset used for evaluation.
The expert prompt shared across all experiments is displayed below.
[…]
3.2.1 Expert vs Amateur 1
In our main method, the amateur is given no context. By doing so, we increase the potential of our context-aware decoding and follow the formulation of (Sanchez et al., 2023)’s method.
The effectiveness on this method is that this will cause it to hopefully rely on the context much more, since the tokens that do rely on it will be much more likelier to occur.
Amateur 1 Prompt A:
3.2.2 Expert vs Amateur 2
In our second amateur, we feed an 8-shot CoT prompt with the questions omitted. This should encourage faithfulness to the question context, as demonstrated in factuality-based tasks by (Shi et al., 2023). That is because it would reduce the probability of logits that would happen even if there’s no question, and increase the probabilities of those that only appear because of it, making the logits that eventually end in it more likely to rely on the question.
[…]
3.2.3 Expert vs Amateur 3
In the third amateur, we feed an 8-shot prompt with questions and answers, but without CoT reasoning. This should hopefully induce further step-by-step reasoning, under the intuition that a reasoning chain uniquely preferred by a CoT prompted model will exhibit more thorough reasoning.
[…]
3.3 Hyperparameter Search
Measuring performance on Mistral-7B (Jiang et al., 2023) using the CommonSenseQA dataset (Talmor et al., 2019), we found an optimal $\alpha$ value of $0.8$ (see  Table 1). This is a constant we used across all datasets and models. In addition, we decided to use 0.5, because although it did not have the second highest accuracy, we wanted another alpha away from 0.8. That would allow us to see a more pronounced difference between the alphas.
$\alpha$ Accuracy
0.5 0.500
0.7 0.620
0.8 0.646
0.9 0.59
Table 1: Accuracy of Different CAD levels while Contrasting CoT and Amateur 3 with Mistral on 200 questions of CommonSenseQA
3.4 Baselines
We first run our datasets through 8-shot CoT prompting, without contrasting with any amateur prompt, setting the benchmark accuracy for the rest of the contrasting experiments. We used the same 8-shot COT prompt as we did for our experts.
4 Results
As seen in Table 2, CAD improved AQUA’s accuracy on Phi by a substantial result, and the same for Mistral on CommonSenseQA. The other accuracy improvements were much more modest, with three percent and below in improvement. The AQUA improvement might not show that much importance, however, because even the improved accuracy still mostly stays at the random rate of guessing.
The CoT Contrast works best with Phi-1.5 on AQuA, while it works best with Mistral 7B on CommonsenseQA. Therefore, we do not find that contrastive CoT increases problem solving performance in all language models for all data.
GSM8K notably shows a significant decrease in accuracy on Phi-1.5 using CoT contrast. This could be due to contamination in the Phi-1.5B dataset (Zhang et al., 2024)
There was also contamination in the Mistral model with GMS8K, so that may have affected the results there too(Zhang et al., 2024).
But the results demonstrate that contrastive CoT is successful when addressing common-sense multiple choice questions. The only time that one of the amateurs does not have a greater value than the baseline on multiple choice is GPT-3.5 on AquA. Some improvements are much more modest than others, but all other model-database combinations that are muultiple choice do improve.
The Mistral multiple-choice offers more improvement than Phi’s multiple-choice which could be due to Mistral being a bigger model with more parameters. It having more parameters could allow it to use the contrasting more powerfully.
Table 5 shows an interesting property of the outputs, as it appears the baseline evaluates its expressions with lower accuracy than most of the other experiments, but still results in a higher accuracy. After conducting analysis on model output for each of the experiments, it appears as though, for the experiments, the mathematical calculations may be higher in accuracy in the Amateurs, but the expressions themselves contain incorrect numerical values which may indicate an incorrect translation from the word questions to expressions.
The overall results demonstrate that contrasting with CoT prompting and amateur models is mostly successful when addressing common-sense multiple choice questions.
Model Dataset Baseline Amateur 1 ($\alpha$ = 0.5) Amateur 1 ($\alpha$ = 0.8)
Phi-1.5 GSM8K 34.3 26.9 23.9
Phi-1.5 AQuA 19.3 22.1 25.6
Phi-1.5 CommonSenseQA 24.5 24.1 24.9
Mistral 7B GSM8K 41.3 42.0 34.6
Mistral 7B AQuA 30.3 33.1 29.1
Mistral 7B CommonSenseQA 47.1 48.7 52.0
Table 2: Accuracies (in percentages) of baseline and contrastive CoT with varying $\alpha$. CCoT typically outperforms standard CoT on arithmetic and commonsense reasoning tasks.
Method GMS8K AQuA CommonSense
Baseline 58.7 52.8 61.4
Amateur 1($\alpha$ = 0.8) 53.9 28.7 64.4
Table 3: GPT-3.5 Results on Amateur 1 on 0.8 Alpha. It does not perform as well on GPT, with only one of the CCoT having a better result.
Model Dataset Baseline Amateur 2 ($\alpha$ = 0.5) Amateur 2 ($\alpha$ = 0.8) Amateur 3($\alpha$ = 0.5) Amateur 3($\alpha$ = 0.8)
Phi-1.5 GSM8K 34.3 30.2 26.3 33.4 31.6
Phi-1.5 AQuA 19.3 23.6 21.3 22.0 21.7
Phi-1.5 CommonSenseQA 24.5 23.5 15.6 24.1 24.9
Mistral 7B GSM8K 41.3 38.9 43.2 42.2 41.6
Mistral 7B AQuA 30.3 35.8 33.9 31.5 29.9
Mistral 7B CommonSenseQA 47.1 26.1 48.2 26.1 39.2
Table 4: Results on Other Amateurs. Amateur 2 seems promising because out of the six categories, it has the best result in four of them.
Baseline Amateur 1 (0.5) Amateur 1 (0.8) Amateur 2 (0.5) Amateur 2 (0.5) Amateur 3 (0.5) Amateur 3 (0.5) Coherence Boosting (0.5) Coherence Boosting (0.8)
Mean 6.110 6.088 6.039 6.071 6.037 6.062 6.089 4.386 5.331
Proportion 0.589 0.617 0.626 0.607 0.615 0.608 0.606 0.529 0.542
Table 5: Comparing mean number of sentences per output and the proportion of expressions evaluated correctly for each experiment. Alpha values are in parenthesis.
5 Related Work
This work is inspired by three main ideas: Context-Aware Decoding, Chain of Thought, and prompting methods. Here we detail the most relevant papers in these paths.
The first direction that this paper relates to is prompting. The last few years have given way to major improvements in input part of prompt instructions. This is apparent with the popularity of few-shot prompting by (Brown et al., 2020), automatically learning prompts (Lester et al., 2021), or providing model prompts by describing tasks (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022). This idea leads to our paper’s relation to Chain-of-Thought Prompting.
Chain-of-Thought prompting was a leading discovery with the ability to improve model performance significantly without scaling size. With such an improvement in performance, one of the main focuses of this paper was trying to take the improvement even further. Our main reference would be (Wei et al., 2023) for its groundbreaking discovery and testing of arithmetic and common-sense benchmarks. This paper uses similar benchmarks and prompting for testing chain of thought with context-aware decoding.
Lastly, this paper relates to others that pushed for context-aware decoding and its ability to decrease hallucination in LLMs. In particular, the paper that was referenced was (Shi et al., 2023) for its novelty in getting LLMs to pay more attention to the context provided. The context-aware decoding was effectively used in this paper to improve the accuracy of models and decrease hallucination.
6 Conclusion
Our findings highlight the impact of integrating expert and amateur models in Context-Aware Decoding across diverse datasets. This study highlights the need for future research prioritizing context-sensitive applications in language modeling. Further, there is a pressing need to explore the scalability of these techniques and their relevance across even broader datasets and across differing models and sizes. Future investigations could delve into testing a wider range of alpha values to unveil potentially more substantial results, aiding our understanding of the nuances within context-driven language modeling.
7 Limitations
Our analysis could be limited in that the difference in strength between the expert and amateur models is too inconsequential. That means that there is a possibility that the models themselves have generated outputs that are not dissimilar enough to impact results.
Another limitation specifically on the Phi 1.5 GMS8K results could be the fact that Phi 1.5 is possibly contaminated with GSM8K (Zhang et al., 2024). This explains why our trials for Phi - 1.5 produced many unfinished and illogical answers when responding to input prompts, yet performed much better on the baseline compared on Amateur 1 on GSM8K (Table 2).
Phi-1.5 has also been known to work best with prompts in a QA format which inherently makes the format of the Amateur 1 and 2 prompts difficult for the model to synthesize outputs, and thus impact results (Li et al., 2023b). This could also result because of the fact that Phi-1.5 is too weak of a model with only 1.3 billion parameters to accurately display the effect of CoT contrasting.
Additionally, GPT-3.5 Turbo is a much larger model compared to Phi-1.5 and Mistral 7B. This could affect results too, because while it seems that the bigger size could make it more effective, as it’s able to do contrasting more effectively, which it might, the contrasting could also get diluted in the billions of parameters.
Among the three, only GPT-3.5 Turbo is an instruction-tuned model, meaning that it’s ability to reason through questions may not be a result of CoT contrasting but rather because it already contains such capabilities. There also might be things similar to our contrasting methods already hidden in their instructions, which would make our methods less effective.
The under-performance of GPT-3.5’s GSM8K and AQuA results may be attributed to GPT-3.5’s difficulty in consistently answering mathematical questions accurately Frieder et al. (2023), as tested GPT versions of January 9, January 30, and GPT-4 extensively with GHOSTS, a natural language mathematics dataset, and concluded they were not competent in advanced mathematical comprehension. This could explain why the accuracy did not improve even with context-aware decoding.
For future experiments, the first steps to improve should be to run experiments on non-instruction tuned and much larger models that can better reflect whether or not chain-of-thought vs. amateur contrasting improves the accuracy of outputs. There could also be more experiments done on different datasets, to see how the constrasting works differently for different types of problems. Additionally, selecting an optimal $\alpha$ value for each individual task, question or even token could potentially improve performance.